OpenStudy (anonymous):

\[Evaluate \int\limits \frac{\sin3x}{\cos ^{5}3x}\]

OpenStudy (anonymous):

u = cos 3x

OpenStudy (anonymous):

see what happens

OpenStudy (anonymous):

\[-\int\limits \frac{du}{u ^{5}}\]

OpenStudy (anonymous):

is it correct?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

missing a 3

OpenStudy (anonymous):

what's next?

OpenStudy (anonymous):

yup typo error.

OpenStudy (anonymous):

missing a 3???

OpenStudy (anonymous):

\[\large \int\limits u^n du=\frac{u^{n+1}}{n+1}+C\]

OpenStudy (anonymous):

yes u missed 3 because du =-3 sin 3x dx

OpenStudy (anonymous):

after letting u=cos 3x u must have \[\int\limits -\frac{1}{3 u^5} du\]

OpenStudy (anonymous):

okay thanks. i get it. :)

OpenStudy (anonymous):

yw :D

OpenStudy (anonymous):

@mukushla what is integ of \[\int\limits -\frac{1}{3u ^{5}}\]

OpenStudy (anonymous):

see my hint up there ; here is n=-5 try to work it out and im here if u have any question

OpenStudy (anonymous):

thank you so much. :)

OpenStudy (anonymous):

yw my friend